Hi,
I'm currently taking Intro to Statistics 101 and I've been baffled by a few problems,which have eluded me for the most of the week. I would be most appreciative of any help with solving the problems below. Thank you for all of your time and help.
Sincerely,
Ken
Type Rh+ 39 35 8 4
Rh- 6 5 2 1 If 2 of the 100 subjects randomly selected
the probability that they are both group O and
type Rh+.

a.assume the selections are made with replacement
b.assume the selections are made without
replacement.

People with blood O and type Rh- are considered
to be universal donors, because they can give
blood to anyone If 4 of the 100 subjects
are randomly selected, find the probability that
they are all universal donors.
a. assume that the selections are made with
replacement.

b. assume that the selections are made without
replacement.

Answer by stanbon(75887)

Group
O A B AB
Type Rh+ 39 35 8 4
Rh- 6 5 2 1
If 2 of the 100 subjects randomly selected
the probability that they are both group O and
type Rh+.
P(a person is O and type Rh+) = 39/100
-------
a.assume the selections are made with replacement
P(2 people are O and Rh+) = 0.39^2 = 0.1521
------------------
b.assume the selections are made without replacement.
(2 people are O and Rh+) = (39/100)(38/99) = 0.1497
==============================
People with blood O and type Rh- are considered
to be universal donors, because they can give
blood to anyone If 4 of the 100 subjects
are randomly selected, find the probability that
they are all universal donors.
a. assume that the selections are made with replacement.
0.39^4
--------------
b. assume that the selections are made without replacement.
39C4/100C4 = 0.0210
=========================
Cheers,
Stan H.

=====================================================

Where does type O fit into the scheme?
What are the numbers for type O?
----------------
Cheers,
Stan H.
==========